OpenStudy (amistre64):

does the sum of a sequence of rational numbers produce a rational number, or not a rational number?

OpenStudy (anonymous):

I'm pretty sure it's rational, but I'm not sure how I'd prove it. Maybe by induction? The sum of the sequence \(\left\{\dfrac{1}{2},\dfrac{1}{4}\right\}\) is rational, since \(\dfrac{1}{2}+\dfrac{1}{4}=\dfrac{3}{4}\). So if you keep adding rational terms I don't see how it could suddenly become irrational.

OpenStudy (amistre64):

i was thinking more along the lines of the zeta function, but then myquesiton is prolly more about the limit of the sum as opposed to a definite value

OpenStudy (amistre64):

\[\Large \zeta(s)=\sum_{n=1}^{\infty}\frac{1}{n^s}\]

OpenStudy (anonymous):

A sequence of rational numbers does not necessarily have to add to give a rational number. A finite sum of rationals will always add to give a rational of course, which can be proved by induction as suggested. But sequences are by definition infinite. The sequence \(\frac{1}{n^2}\) for instance, sums to give \(\frac{\pi^2}{6}\). (Not sure I can show that to be honest....)

OpenStudy (amistre64):

euler did a nice job of showing it :)

OpenStudy (anonymous):

Isn't \(s\) a complex number?

OpenStudy (anonymous):

Strictly speaking, a real number is a complex number...

OpenStudy (amistre64):

s can be defined as a complex number, but can be rational

OpenStudy (amistre64):

1/n^2 sums to pi^2/6 as Erin mentions

OpenStudy (amistre64):

http://www.youtube.com/watch?v=wcjknbmTYOI this gives a nice rundown of the "showing"

OpenStudy (amistre64):

around 35 to 37 gets into the meat of it .... unless you want to watch it from the beginning :)

OpenStudy (anonymous):

Oooh great :) We were told we'd see a proof of it when we study Fourier series next year, but an elementary proof would be nice.

OpenStudy (amistre64):

i keep trying to learn about foureir series and keep falling asleep

OpenStudy (amistre64):

having a pretty teacher might hold my interest, but then i run the risk of being distracted too much to learn :)

OpenStudy (anonymous):

With that introduction, I'll look forward to Fourier series next term then ;-) Just found this proof: http://www.maths.usyd.edu.au/u/daners/publ/abstracts/zeta2/zeta2.pdf Rather neat!